# Velocity Jumps for Molecular Dynamics

### v1

@@ 1 Introduction
The paper is organized as follows. Section 2 is devoted to the description of the numerical method: we define in Sections 2.1 and 2.2 the continuous-time hybrid process, Sections 2.3 and 2.4 describe the simulation techniques and how they yield a computational advantage. Then, in Section 3, we see how our approach can be applied to Coulomb and van der Waals interactions in a classical force field. Finally, Section 4 is devoted to numerical experiments, both on CPU and GPU versions of the Tinker-HP software [29]. The Appendix 6 gives the proof of some of the results and other technical details.
@@ 2.1 The velocity jump Langevin process
with the velocity $V_{t}$ undergoing additional random jumps at rate $\lambda_{i}(x,v)$ following a jump kernel $q_{i}(x,v,\text{d}v^{\prime})$ (that describes the probability distribution of the velocity after a jump), both defined below in Section 2.2, as also explained in [48] in a particular case. The jump mechanism $(\lambda_{i},q_{i})$ depends on the force $F_{i}$ in a way that ensures that the equilibrium measure of the process in indeed the canonical measure (1). Informally, the process follows (3) but, between times $t$ and $t+\delta$ for a small $\delta$, with probability $\lambda_{i}(X_{t},V_{t})\delta+o(\delta)$, its velocity $v$ is re-sampled according to the kernel $q_{i}(X_{t},V_{t},\text{d}v^{\prime})$. It can be seen as a hybrid between a diffusion and a piecewise deterministic Markov process (PDMP) such as considered in [51, 52, 53], which have recently drawn much attention in numerical probabilities and Bayesian statistics, see [52, 54, 55, 56, 57, 58]. As we will explain, simulating this process in a certain way yields a computational advantage over the simulation of the Langevin diffusion, because the computation of the force $F_{i}$ will only be required when a jump of type $i$ is proposed, which will not happen at each time step for every $i$. We refer to the Appendix 6.1 for a more rigorous definition.
@@ 2.2 The jump mechanism
Moreover, as explained in Appendix 6.4, the definitions (6) and (7) ensure that the Gibbs measure $\mu$ is left invariant by the velocity jump Langevin process defined in Section 2.1, and the ergodicity of the process in a theoretical setting is proven in [49]. In other words, when $\varepsilon$ is small enough, the velocity jump Langevin process is close, in terms of stochastic trajectories, to the Langevin dynamics, which yields (as we will see in Section 4), a good preservation of dynamical properties of the system on top of a computational speedup without loss of accuracy in sampling.
@@ 2.3.1 JUMP integrator
In practice, since the solution of the SDE (3) cannot be simulated exactly, time discretization methods need to be used to approximate the process. Similarly to the BAOAB [38, 39, 40] scheme, the continuous-time dynamics of the velocity jump Langevin process can be approximated following a Trotter/Strang splitting scheme. We will describe three of them, and refer to the Appendix 6.2 for more technical details.
@@ 2.3.2 JUMP-RESPA and JUMP-RESPA1 integrators
This procedure can itself be iterated to include more different time steps, such as BAOAB-RESPA1 [34] that has three layers of time steps. By analogy, We will refer to this situation as the JUMP-RESPA1 integrator. More details are given in Appendix 6.2.
@@ 2.4 Efficient simulation of jumps
Let us start with the process $(\widetilde{W}_{s})_{s\geqslant 0}$ (corresponding to the (C) step in the CAOAC scheme). We suppose that each jump rate $\lambda_{i}$ is bounded by a constant $\overline{\lambda_{i}}$, and denote $\overline{\lambda}=\sum_{i}\overline{\lambda_{i}}$. Then an easy computation on the generator (see Appendix 6.3) shows that the jumps can be simulated exactly this way: starting from $(x,v)$,
@@
In fact, as shown in Appendix 6.6, in the general case, the bounds $\overline{\lambda}_{i}$ are local (they depend the velocity $v$) and need to be re-evaluated after each jump. This is not a problem, since the only requirement in order to have a computational advantage is that those bounds do not depend on $F_{i}$. However, in the case of the Bouncy Particle Sampler (i.e. the jump mechanism defined in (4) and (5)), the norm of the velocity doesn’t change after a jump, so $\overline{\lambda}$ stays constant throughout the time step. In those cases, when we simulate the process up to a time $\delta$, we don’t need to know exactly the jump times, but only the number of jumps and the order in which they occur. Classical properties of the exponential distribution ensure that during the time interval $\left[0,\delta\right]$, the number of jump proposals (step 1 above) follows a Poisson distribution of parameter $\overline{\lambda}\delta$. Then the simulation procedure for the (J) part of BJAOAJB is the following: starting from $(x,v)$,
@@ 3.1 Direct electrostatic jumps
As reminded in the Appendix 6.5, the Ewald summation splits the electrostatic interactions into a direct and a reciprocal part. In Periodic Boundary Conditions (PBC), this decomposition can be chosen in a way that enforces the “minimum image convention”: in the direct space, each atom interacts only with the closest image of an other atom. We assume in the following that this condition is satisfied. By denoting $r_{ij}=|x_{i}-x_{j}|$ the distance between the atoms $i$ and $j$, the direct electrostatic energy associated to each pair $(i,j)$ is:
@@
In order to decompose this term into a short and a long range contribution, let $r_{c}$ be a cutoff (smaller than the “Ewald cutoff” that separates the direct and reciprocal part, whose value is enforced by the choice of the parameter $\alpha$), $\chi(r)$ a switching function that goes smoothly from 1 to 0 around the cutoff (its precise formula is given in Appendix 6.6) and let
@@
where $\widetilde{G}$ has been generated from the distribution $f_{\eta}$ with $\eta=\varepsilon(x)\sqrt{\beta m_{i}}\frac{v_{i}\cdot F_{ij}(x)}{|F_{ij}(x)|}$. As detailed in the Appendix 6.6, in this case it is easy to obtain explicit analytic bounds on $|F_{ij}|$ that are required in the thinning method. Moreover, for each $i$, those bounds are uniform in $j$ which, as explained in Section 2.4, simplifies the thinning procedure.
@@
7:         $a_{i}\leftarrow-\nabla U_{short}(x)/mass(i)$triangleright a is the acceleration due to the short range forces
@@ 4 Numerical experiments
All variations of the algorithms described in Section 3 have been implemented in Tinker-HP software for molecular dynamics, both on the CPU and the parallel GPU versions. The software will be made freely available to academic users within the next release of the Tinker-HP code [65, 66, 21, 29]. Computations were done on the computing cluster of the Theoretical Chemistry Laboratory (LCT) of Sorbonne University, Paris, on a 22 cores Intel(R) Xeon(R) Silver 4116 CPU @ 2.10GHz for the CPU version, and a Nvidia Quadro GV100 for the GPU version. All numerical tests were done on water systems of sizes ranging from 216 to 96000 molecules, with SPC/Fw model in the OPLS-AA force-field, in periodic boundary conditions. More details on the implementation and the simulation parameters are given in the Appendix 6.7 and 6.7.2.
@@
Those differences between GPU and CPU have different reasons. First, the neighbor lists are treated differently, see the Appendix 6.7.2 for more details. Then, the proportion of time spent in the different routines of the integrator is different on CPU and GPU. For instance, the parallel architecture of GPUs allows for a more efficient treatment of the van der Waals and direct electrostatic interactions, while the many-body Ewald sum takes almost $50\%$ of the total computation time, especially on larger systems. As a consequence, on a large system, reducing the reciprocal part of electrostatics and treating a larger portion of the potential with jumps yields a significant computational advantage. For reference, Table 3 gives the proportion of time spent in each of the main routines of the CPU and GPU implementations, comparing two system sizes.

### v2

@@ 1 Introduction
The paper is organized as follows. Section 2 is devoted to the description of the numerical method: we define in Sections 2.1 and 2.2 the continuous-time hybrid process, Sections 2.3 and 2.4 describe the simulation techniques and how they yield a computational advantage. Then, in Section 3, we see how our approach can be applied to Coulomb and van der Waals interactions in a classical force field. Finally, Section 4 is devoted to numerical experiments, both on CPU and GPU versions of the Tinker-HP software [29]. The supporting information gives the proof of some of the results and other technical details.
@@ 2.1 The velocity jump Langevin process
with the velocity $V_{t}$ undergoing additional random jumps at rate $\lambda_{i}(x,v)$ following a jump kernel $q_{i}(x,v,\text{d}v^{\prime})$ (that describes the probability distribution of the velocity after a jump), both defined below in Section 2.2, as also explained in [48] in a particular case. The jump mechanism $(\lambda_{i},q_{i})$ depends on the force $F_{i}$ in a way that ensures that the equilibrium measure of the process in indeed the canonical measure (1). Informally, the process follows (3) but, between times $t$ and $t+\delta$ for a small $\delta$, with probability $\lambda_{i}(X_{t},V_{t})\delta+o(\delta)$, its velocity $v$ is re-sampled according to the kernel $q_{i}(X_{t},V_{t},\text{d}v^{\prime})$. It can be seen as a hybrid between a diffusion and a piecewise deterministic Markov process (PDMP) such as considered in [51, 52, 53], which have recently drawn much attention in numerical probabilities and Bayesian statistics, see [52, 54, 55, 56, 57, 58]. As we will explain, simulating this process in a certain way yields a computational advantage over the simulation of the Langevin diffusion, because the computation of the force $F_{i}$ will only be required when a jump of type $i$ is proposed, which will not happen at each time step for every $i$. We refer to the SI, Section 1, for a more rigorous definition.
@@ 2.2 The jump mechanism
Moreover, as explained in SI, section 4, the definitions (6) and (7) ensure that the Gibbs measure $\mu$ is left invariant by the velocity jump Langevin process defined in Section 2.1, and the ergodicity of the process in a theoretical setting is proven in [49]. In other words, when $\varepsilon$ is small enough, the velocity jump Langevin process is close, in terms of stochastic trajectories, to the Langevin dynamics, which yields (as we will see in Section 4), a good preservation of dynamical properties of the system on top of a computational speedup without loss of accuracy in sampling.
@@ 2.3.1 JUMP integrator
In practice, since the solution of the SDE (3) cannot be simulated exactly, time discretization methods need to be used to approximate the process. Similarly to the BAOAB [38, 39, 40] scheme, the continuous-time dynamics of the velocity jump Langevin process can be approximated following a Trotter/Strang splitting scheme. We will describe three of them, and refer to the SI, section 2, for more technical details.
@@ 2.3.2 JUMP-RESPA and JUMP-RESPA1 integrators
This procedure can itself be iterated to include more different time steps, such as BAOAB-RESPA1 [34] that has three layers of time steps. By analogy, We will refer to this situation as the JUMP-RESPA1 integrator. More details are given in SI, section 2.
@@ 2.4 Efficient simulation of jumps
Let us start with the process $(\widetilde{W}_{s})_{s\geqslant 0}$ (corresponding to the (C) step in the CAOAC scheme). We suppose that each jump rate $\lambda_{i}$ is bounded by a constant $\overline{\lambda_{i}}$, and denote $\overline{\lambda}=\sum_{i}\overline{\lambda_{i}}$. Then an easy computation on the generator (see SI, section 3) shows that the jumps can be simulated exactly this way: starting from $(x,v)$,
@@
In fact, as shown in SI, section 6, in the general case, the bounds $\overline{\lambda}_{i}$ are local (they depend the velocity $v$) and need to be re-evaluated after each jump. This is not a problem, since the only requirement in order to have a computational advantage is that those bounds do not depend on $F_{i}$. However, in the case of the Bouncy Particle Sampler (i.e. the jump mechanism defined in (4) and (5)), the norm of the velocity doesn’t change after a jump, so $\overline{\lambda}$ stays constant throughout the time step. In those cases, when we simulate the process up to a time $\delta$, we don’t need to know exactly the jump times, but only the number of jumps and the order in which they occur. Classical properties of the exponential distribution ensure that during the time interval $\left[0,\delta\right]$, the number of jump proposals (step 1 above) follows a Poisson distribution of parameter $\overline{\lambda}\delta$. Then the simulation procedure for the (J) part of BJAOAJB is the following: starting from $(x,v)$,
@@ 3.1 Direct electrostatic jumps
As reminded in the SI, section, 5, the Ewald summation splits the electrostatic interactions into a direct and a reciprocal part. In Periodic Boundary Conditions (PBC), this decomposition can be chosen in a way that enforces the “minimum image convention”: in the direct space, each atom interacts only with the closest image of an other atom. We assume in the following that this condition is satisfied. By denoting $r_{ij}=|x_{i}-x_{j}|$ the distance between the atoms $i$ and $j$, the direct electrostatic energy associated to each pair $(i,j)$ is:
@@
In order to decompose this term into a short and a long range contribution, let $r_{c}$ be a cutoff (smaller than the “Ewald cutoff” that separates the direct and reciprocal part, whose value is enforced by the choice of the parameter $\alpha$), $\chi(r)$ a switching function that goes smoothly from 1 to 0 around the cutoff (its precise formula is given in SI, section 6) and let
@@
where $\widetilde{G}$ has been generated from the distribution $f_{\eta}$ with $\eta=\varepsilon(x)\sqrt{\beta m_{i}}\frac{v_{i}\cdot F_{ij}(x)}{|F_{ij}(x)|}$. As detailed in the SI, section 6, in this case it is easy to obtain explicit analytic bounds on $|F_{ij}|$ that are required in the thinning method. Moreover, for each $i$, those bounds are uniform in $j$ which, as explained in Section 2.4, simplifies the thinning procedure.
@@
7:         $a_{i}\leftarrow-\nabla U_{\text{short}}(x)/m_{i}$triangleright a is the acceleration due to the short range forces
@@ 4 Numerical experiments
All variations of the algorithms described in Section 3 have been implemented in Tinker-HP software for molecular dynamics, both on the CPU and the parallel GPU versions. The software will be made freely available to academic users within the next release of the Tinker-HP code [65, 66, 21, 29]. Computations were done on the computing cluster of the Theoretical Chemistry Laboratory (LCT) of Sorbonne University, Paris, on a 22 cores Intel(R) Xeon(R) Silver 4116 CPU @ 2.10GHz for the CPU version, and a Nvidia Quadro GV100 for the GPU version. All numerical tests were done on water systems in the NVT ensemble with temperature $T=300K$, of sizes ranging from 216 to 96000 molecules, with SPC/Fw model in the OPLS-AA force-field, in periodic boundary conditions. Simulation parameters are given at the end of this section, and SI, section 7, gives more details on the implementation.
@@
Those differences between GPU and CPU have different reasons. First, the neighbor lists are treated differently, see the end of this section for more details. Then, the proportion of time spent in the different routines of the integrator is different on CPU and GPU. For instance, the parallel architecture of GPUs allows for a more efficient treatment of the van der Waals and direct electrostatic interactions, while the many-body Ewald sum takes almost $50\%$ of the total computation time, especially on larger systems. As a consequence, on a large system, reducing the reciprocal part of electrostatics and treating a larger portion of the potential with jumps yields a significant computational advantage. For reference, Table 3 gives the proportion of time spent in each of the main routines of the CPU and GPU implementations, comparing two system sizes.